OpenStudy (anonymous):

given f(x)=x^2+3 and g(x)=x+5/x. Find (g o f)(4) [Hint: (g o f)(4)=g(f(4))]

OpenStudy (welshfella):

first find f(4) you do this by replacing the x in the expression for f(x) by 4 and working it out. can you do that?

OpenStudy (welshfella):

f(x) = x^2 + 3 f(4) = 4^2 + 3 = ?

OpenStudy (anonymous):

19